# Ageing-aware Energy Management for Residential Multi-Carrier Energy Systems

### v3

@@ 5.3 Case Study III: Manageing aged and fresh batteries
Figure 15 presents the degradation patterns for the 2 planners and different $SoH$. In the aged battery, the share of calendar ageing (against the total) within $i_{\text{SEI},b,t}$ is much smaller and thus the percentual $Q_{\textrm{loss}}$ is almost 35% smaller than in the new battery packs. In both seasons, the EV s patterns are similar for both BNoDeg and CPBDeg. The change in EV trajectories between aged and fresh cells shows a slight increase $FEC$, due to a smaller rated capacity, and a decrease in relative ageing, due to reduced calendar ageing. The impact on the BESS is more pronounced. In winter, CPBDeg has a smaller $\frac{\partial Q_{\textrm{loss}}}{\partial FEC}$ (upper-right hand side of the graph) for both aged and new cells, with BNoDeg doing more eq. cycles. In the winter CPBDeg still has a smaller $\frac{\partial Q_{\textrm{loss}}}{\partial FEC}$ in the new cell but not in the used cell, where this is achieved by the benchmark controller. In other words, as the battery ages, the predominance of calendar ageing within the SEI layer is diminished. Thus, the link between $i_{b,t}$ and $Q_{\text{loss},b,t}$ becomes stronger, due to the predominance of active material loss and the cyclic component of the SEI layer.
@@
Finally, when adding the costs to the analysis, summarized in Table 3, it is noticed that CPBDeg achieves lower capacity fade $Q_{\textrm{loss}}$ than BNoDeg across all seasons and $SoH$. Even improving the total grid cost $C_{\textrm{grid}}$ as the cells degrade. In particular during summer, with a 35% observed improvement. On the other hand, BNoDeg worsens its performance as the cells degrade with higher $C_{\textrm{grid}}$, $Q_{\textrm{loss}}$ and model bias.
In summary, the performance of the proposed CPBDeg planner comparatively improves when using used cells, with respect to its linear BNoDeg benchmark. This is because the ageing PBROM s encapsulate the fade of calendar ageing ($\sqrt{t}$) and rise of cyclic SEI and AM ($i_{b,t},SoC_{b,t}$) as the battery ages.
@@
CPBDeg 16.3 78.7 130.0 149.0
BNoDeg 15.2 77.6 132.3 158.3
Table 3: Cost and degradation summary.
@@ 5.4 Limitations & Future works
  • 1.
    This paper assumes that battery pack temperatures are constant (BESS and EV), since temperature control is part of the local primary controls performed by the respective BMS. Future research could explore active thermal controls and their impact on degradation control.
  • 2.
    Even though the terminal set is more flexible than the standard literature, the design is still arbitrary. It would be interesting to design the terminal conditions as part of a wider control scheme, such as seasonal optimization.
  • 3.
    Forecasts are needed for all exogenous processes $W_{t+1}$.
  • 4.
    State observers are also necessary for all devices in an experimental application to correctly feed back the states $S_{a,t}$ to the optimization.
  • 5.
    The thermal flow-based models used in this paper assume a fixed building temperature decided in an external system. This is the assumption behind $P_{\textrm{load}}^{\textrm{th}}$. Flexible thermal controls could also be integrated.
  • 6.
    Uncertainty handling is limited since re-optimization frequency is only 24 hours. Future works include improving this through dynamic programming approaches and/or economic model predictive control.
@@ 6 Conclusions & Discussion
In summary, this paper presents an optimization-based day-ahead planner for residential multi-carrier energy systems that uses PBROM models to integrate battery ageing. Its integration with the thermal and mobility carriers is achieved by designing specific terminal sets for the individual storage systems, depending on their use and technology. The proposed CPBDeg planner can handle different cathode chemistries as well as batteries in different ageing states. The planner does this with lower degradation than the benchmarks. The PBROM integration comes at the expense of slightly higher computational times and grid cost in the case of NMC cells.
@@
In the first case study, it is shown how advanced PBROM can be used to reduce battery ageing while maintaining a grid cost comparable to benchmarks, in accordance with the literature for standalone utility-scale applications [41, 27, 65, 39, 38]. This is because the planner can better control the degradation per full equivalent cycle than the benchmarks.
@@
In the second case study, the proposed planner is equipped with battery packs of different cathode chemistries, and its performance is compared. Empirical models are cathode-specific and thus can’t handle multiple cell chemistries. The LFP battery pack has a lower total $Q_{\textrm{loss}}$ as per established knowledge and can achieve lower grid costs $C_{\textrm{grid}}$ than its NMC counterpart. Even more so, when considering that the relative cost $c_{\textrm{loss}}$ for both packs was the same, when in reality LFP packs have a lower cost than NMC packs. Thus, Section 5.2 is a conservative approximation, and LFP packs have the potential to enable even lower operation costs. The grid savings are achieved just by changing some simple model parameters taken from the literature. This flexibility is an essential feature because it allows the EMS to exploit any battery pack at hand fully.
@@
The last case study showed how the proposed EMA handles aged and new batteries seamlessly, even improving performance as the battery ages, improving the summer grid cost by 35%. The reason is that the EMA can identify predominant degradation mechanisms from the ageing PBROM, exploiting the decrease of SEI importance over time and the rise of active material loss. Its linear counterpart, however, is unaware of the degradation, increasing the number of rejected actions. Summing up, grid costs can be reduced by using LFP battery packs or modifying objective weights, the latter at the expense of higher degradation. Finally, the integration of ageing PBROM s has a low impact on the HP and TESS, meaning that the almost 25% of grid cost savings of Section 5.2 are achieved without fundamentally modifying the thermal strategy. In applications where scalability and fast computation are key drivers, decomposition strategies could be explored to exploit this coupling.
@@ 8 Declaration of generative AI and AI-assisted technologies in the writing process
Appendix A Appendix: A
Graphite anode open-circuit voltage:

### v4

@@ 5.3 Case Study III: Managing aged and fresh batteries
Figure 16 presents the degradation patterns for the 2 planners and different $SoH$. In the aged battery, the share of calendar ageing (against the total) within $i_{\text{SEI},b,t}$ is much smaller and thus the percentual $Q_{\textrm{loss}}$ is almost 25% smaller than in the new battery packs. In both seasons, the EV s patterns are similar for both BNoDeg and CPBDeg. The change in EV trajectories between aged and fresh cells shows a slight increase $FEC$, due to a smaller rated capacity, and a decrease in relative ageing, due to reduced calendar ageing. The impact on the BESS is more pronounced. In winter, CPBDeg has a smaller $\frac{\partial Q_{\textrm{loss}}}{\partial FEC}$ (upper-right hand side of the graph) for both aged and new cells, with BNoDeg doing more eq. cycles. In the winter CPBDeg still has a smaller $\frac{\partial Q_{\textrm{loss}}}{\partial FEC}$ in the new cell but not in the used cell, where this is achieved by the benchmark controller. In other words, as the battery ages, the predominance of calendar ageing within the SEI layer is diminished. Thus, the link between $i_{b,t}$ and $Q_{\text{loss},b,t}$ becomes stronger, due to the predominance of active material loss and the cyclic component of the SEI layer.
@@
Finally, when adding the costs to the analysis, summarized in Table 6, the CPBDeg achieves lower capacity fade $Q_{\textrm{loss}}$ than BNoDeg across all seasons and $SoH$ (between 30-45% less). Even improving the total grid cost $C_{\textrm{grid}}$ by 6% in the summer as the cells degrades. On the other hand, BNoDeg worsens its performance as the cells degrade with higher $C_{\textrm{grid}}$ (6-9%), $Q_{\textrm{loss}}$ and model bias.
In summary, the performance of the proposed CPBDeg eMPC controller comparatively improves when using used cells, with respect to its linear BNoDeg benchmark. This is because the ageing PBROM s encapsulate the fade of calendar ageing ($\sqrt{t}$) and rise of cyclic SEI and AM ($i_{b,t},SoC_{b,t}$) as the battery ages.
@@
BNoDeg 58.66 111.31 281.11 288.31
CPBDeg 60.87 120.86 168.64 149.30
Table 6: Cost and degradation summary with weights $w_{\text{grid}}=1,\ w_{\text{loss}}=0.1,\ w_{\text{SoC}}=w_{\text{TESS}}=1000$.
@@ 5.4 Limitations & Future works
As it was mentioned in Section 3 two major assumptions are made regarding the battery packs: (i) moderate C-rates of less than 1C, (ii) constant comfortable battery pack temperature. Both allow us to neglect important factors, such as Li-plating. Temperature control is part of the local primary controls performed by the respective BMS and no-fast charging outside of the house assumed. Of course, they limit the extension of the presented results to specific days of high heat or low temperatures, since other degradation mechanisms will play a more relevant role. Moreover, as the operation bounds grow more extreme (C-rate and temperatures), the less accurate the models grow even PB models. Future research could explore active thermal controls and their impact on the presented degradation control.
The thermal flow-based models used in this paper assume a fixed building temperature decided externally, that produces the exogenous thermal load $P_{\textrm{load}}^{\textrm{th}}$. Another thermal limitation is that efficiencies and conversion factors are assumed linear. These limits the presented work thermal flexibility, since the building’s thermal capacity could be used as a passive thermal storage and HP consumption prediction could be improved. Both are explored in other works such as [Slaifstein2026SequentialMPC, Alpizar-Castillo2024ModellingHouse].
A natural limitation of all model-based optimal control methods is that they rely on accurate uncertainty estimation techniques. This means that accurate forecasters are needed to feed belief states $\tilde{B}_{t}$ of inputs $W_{t+1}$ and state observers/estimators to feedback state measurements $S_{a,t+1}$ and accurately parametrize the transition functions $\tilde{S}^{M}_{t}(\cdot)$. The accuracy of both the forecasts $\tilde{B}_{t}$ and device models $\tilde{S}^{M}_{a,t}(\cdot)$ is a necessary condition for successful real-life policies. The development of battery observers is outside of the scope of this paper but the reader might refer to [Fan2023NondestructivePaths, Li2024NonlinearCharging, Plett2024BatteryMethods, Couto2025DegradationApproach]. The experimental validation of particular models used is outside of the scope of this paper, since we do not want to contribute with a new battery model but with a new dispatch algorithm. The core idea of the present paper is showing that ageing PBROM s, such as the one developed by Jin [Jin2022] or similar, can be integrated to wider EMS/control schemes.
Regarding the periodic constraints of the BESS and TESS. Even though the terminal set of the BESS is more flexible than the standard literature, the design is still arbitrary. It would be interesting to design the terminal conditions as part of a wider control scheme, such as seasonal optimization [Darivianakis2017AManagement]. Not only for the BESS but for the TESS, since due to the $H^{\text{DA}}=48$hs, $H^{\text{MPC}}=24$hs its low short-term efficiency leads to discharging and not cycling the TESS.
The practical implications for residential consumers are twofold: the short-term grid-cost savings often come at the expense of accelerated battery aging. While the BNoDeg and CEmpDeg controller can yield lower monthly grid costs ($C_{\text{grid}}$), they do so by aggressively cycling the battery in ways that ignore internal degradation mechanisms. By utilizing the CPBDeg control, users can transition from a short-sighted “bill minimization” strategy to a “Total Cost of Ownership” (TCO) approach. Given the capacity fade cost of approximately 300 €/kWh, the ability of the PBROM-based EMS to reduce degradation per cycle ($\partial Q_{\text{loss}}/\partial FEC$) suggests that the battery’s operational lifespan could be extended by several years. This significantly improves the Net Present Value (NPV) of the residential energy system and reduces the frequency of capital-intensive storage replacements. Naturally, further work is needed to implement similar schemes in real life and until then this work acts as a benchmark to highlight potential savings or practical recommendations. These include, choosing LFP battery packs or maintaining batteries close to their $\underline{SoC}_{b,t}$.
@@ 6 Conclusions & Discussion
In summary, this paper presents an optimization-based non-linear eMPC for residential multi-carrier energy systems that uses PBROM models to integrate battery ageing. Its integration with the thermal and mobility carriers is achieved by designing specific terminal sets for the individual storage systems, depending on their use and technology. The proposed CPBDeg control can handle different cathode chemistries as well as batteries in different ageing states. The controller can do this with lower degradation than the benchmarks, by increasing $w_{\text{loss}}$. The PBROM integration comes at the expense of slightly higher computational times and grid cost in the case of NMC cells.
@@
In the first case study, it is shown how advanced PBROM can be used to reduce battery ageing while maintaining a grid cost comparable to benchmarks, in accordance with the literature for standalone utility-scale applications [Reniers2018, Reniers2019, Reniers2020, Reniers2021, Reniers2023]. Moreover, the CPBDeg has the most smallest capacity fade per cycle. This is because the control can effectively relate actions to degradation states.
@@
In the second case study, the proposed planner is equipped with battery packs of different cathode chemistries, and its performance is compared. Empirical models are cathode-specific and thus can’t handle multiple cell chemistries. The LFP battery pack has a lower total $Q_{\textrm{loss}}$ as per established knowledge and can achieve lower grid costs $C_{\textrm{grid}}$ than its NMC counterpart. Even more so, when considering that the relative cost $c_{\textrm{loss}}$ for both packs was the same, when in reality LFP packs have a lower cost than NMC packs. Thus, Section 5.2 is a conservative approximation, and LFP packs have the potential to enable even lower TCO. The grid savings are achieved just by changing some simple model parameters taken from the literature. This flexibility is an essential feature because it allows the EMS to exploit any battery pack at hand fully. This directly impacts the choice of the residential users towards LFP battery packs, because even without aging-aware controls the overall ageing favors LFP packs which is already an industry standard.
@@
The last case study showed how the proposed EMA handles aged and new batteries seamlessly, even improving performance as the battery ages, improving the summer grid cost by 6%. The reason is that the EMA can identify predominant degradation mechanisms from the ageing PBROM, exploiting the decrease of SEI importance over time and the rise of active material loss. Its linear counterpart, however, is unaware of the degradation causing the energy costs to increase. Summing up, grid costs can be reduced by using LFP battery packs or modifying objective weights, the latter at the expense of higher degradation. Finally, the integration of ageing PBROM s has a low impact on the HP and TESS, meaning that the grid cost savings during summer of Section 5.2 are achieved without fundamentally modifying the thermal strategy. In applications where scalability and fast computation are key drivers, decomposition strategies could be explored to exploit this coupling.
On a broader scope, even though relevant challenges are still pending or out of the scope of this paper, this paper aims to show the potential savings enabled by aging-aware control laws in the residential sector. The findings presented here can be easily transferred to other sectors such as commercial or industry, which might have more tools to implement complex EMS.
For physical setups the proposed physics-based approach requires a non-linear observer to identify internal states and a system identification algorithm to parametrize the models. Adaptive control techniques/ online learning techniques are crucial for scaling implementation. For the case of the BESS initial tests and model identification can be done offline before start-up, and even offline during operation if historic data is continuously stored. However, for EV s models, parametrization deems a challenge since a previously unknown car may appear or due to unknown driving conditions and profiles. Thus, the development of effective and accurate observers to identify and parametrize PBROM online automatically is crucial.
@@ 8 Declaration of generative AI and AI-assisted technologies in the writing process
Appendix A Building Battery Models
Graphite anode open-circuit voltage [Jin2017, Jin2022]:
@@ Appendix A Building Battery Models
Figure 17: ECM System Identification pipeline used to build planner models.
In order to test the each of the EMS decision-making a virtual simulator is necessary to stablish a common ground to compare policies [Powell2022]. A reasonable tool is the PyBaMM modelling and simulation library as a digital simulator, which is widely used in the battery community for design, optimization and benchmarking [Sulzer2021PythonPyBaMM]. PyBaMM has datasets for several real cells, including NMC and LFP cells [Okane2022, Prada2013Simulations].
To construct the models used in the optimizer, $\tilde{S}^{M}_{b,t}(\cdot)$, and avoid using physical cells, one can use a PyBaMM model (compendium of PDEs) and simulate dynamic and open circuit tests as the ones specified in [Plett2016], and identify an ECM using system identification with the ESCtoolbox, again following [Plett2016]. The pipeline is presented in Figure 17.
To build the simulator, $S_{b,t}^{M}(\cdot)$, directly using PyBaMM is inconvenient because interfacing Julia and Python is possible but hinders execution time. The solution is using the LiiBRA.jl package [Planden2022], which builds a state-space PBROM from the SPM or DFN model using Discrete-time Realization Algorithms (DRA). This greatly improves simulation times and allows for detailed simulations with a time resolution of $\Delta t_{s}=1$s [Planden2022].
Appendix B Periodic Condition
In Eq. LABEL:eq:besspcond the value of $t_{1}=6$hs was chosen experimentally through simulations. Figure 18 presents the $SoC_{\text{BESS},t}$ and the total objective $J$ as function of terminal condition $t_{1}$ for the standard (mean inputs) weeks of summer and winter. In it, the $t_{1}=6$hs maintains the second place in both weeks, with its closest competitor of $t_{1}=12$hs is dependent of solar generation which plummets in winter.
(a) Summer
(b) Winter
Figure 18: BESS State-of-Charge $SoC_{\text{BESS},t}$ and total objective $J$ as a function of terminal condition $t_{1}$ in (a) summer and (b) winter.
Glossary
Acronyms
AM
active material loss
BESS
battery energy storage system
BM
bucket model
BMS
battery management system
DER
distributed energy resources
DLA
Direct Lookahead
ECM
equivalent circuit model
EMA
energy management algorithm
eMPC
economic Model Predictive Control
EMS
energy management systems
ESS
energy storage systems
EV
electric vehicle
HESS
hybrid energy storage system
HP
heat pump
LFP
Lithium iron phosphate
MCES
multicarrier energy systems
MPC
Model Predictive Control
NMC
nickel manganese cobalt oxides
OCP
optimal control problem
PB
physics-based
PBROM
physics-based reduced order model
PEI
power electronic interface
RL
reinforcement learning
SCM
series connected module
SPD
sequential decision problem
SEI
solid electrolyte interface
SPV
solar photovoltaics
ST
solar thermal
TESS
thermal energy storage system
UMF
Universal Modeling Framework